OpenStudy (anonymous):

The line y − 6 = 2(x + 4) passes through which one of the following points? a.(-4,-6) b.(4,-6) c.(-4,6) d.(2,6)

OpenStudy (notamathgenius):

Just plug in the numbers for X and Y

OpenStudy (anonymous):

i don't know how

OpenStudy (notamathgenius):

Ok you see Y right? well take A for example and put -6 for y and -4 for X and see if they equal the same answer

OpenStudy (notamathgenius):

If they don't go down the list till they do

OpenStudy (anonymous):

it comes up as an error

OpenStudy (notamathgenius):

y − 6 = 2(x + 4) -6 - 6 = 2(-4+4) you do PEMDAS parenthesis first

OpenStudy (anonymous):

i did the top and got y=2x+14

OpenStudy (notamathgenius):

Good, now take (-4,-6) again and use it for that

OpenStudy (anonymous):

how woudl i put that

OpenStudy (notamathgenius):

y=2x+14 -6=2(-4)+14

OpenStudy (anonymous):

i do that but its says it a function

OpenStudy (notamathgenius):

2(-4)=-8+14=6 doesn't equal -6 so a is wrong

OpenStudy (anonymous):

im going to go with it being b

OpenStudy (notamathgenius):

You sure?

OpenStudy (anonymous):

not really

OpenStudy (notamathgenius):

Well, do what we did with A over again it's better than getting it wrong, correct?

OpenStudy (anonymous):

yup

OpenStudy (notamathgenius):

Ok, just do it again

OpenStudy (anonymous):

my friend said its c

OpenStudy (notamathgenius):

It is c 0=0

OpenStudy (anonymous):

sweet